OpenStudy (anonymous):

How can x2 = x2 + 2x + 9 be set up as a system of equations? y = x2 - 9 y = x2 + 2x + 9 y = x2 y = x2 + 2x + 9 y = x2 + 2x y = x2 + 9 y = x2 y = 2x + 9

OpenStudy (anonymous):

@ikram002p

OpenStudy (ikram002p):

what do u think ?

OpenStudy (anonymous):

the first one, but i dont feel like im right

OpenStudy (ikram002p):

yeah its not right , ill tell u how to make sure

OpenStudy (ikram002p):

\(x^2 = x^2 + 2x + 9\) so take each side as equation it self

OpenStudy (ikram002p):

\(y= x^2 + 2x + 9=x^2\) so \(y= x^2 + 2x + 9 \) \(y= x^2\)

OpenStudy (anonymous):

allright, so its either the second or last

OpenStudy (mathmale):

Note that you now have two equations that begin with "y=", and as such you have a set of simultaneous equations (a system of equations) that are to be solved for both x and y. To answer your question: NO. It's not a matter of "either...or" here. You still need both equations to find a solution. A solution, if it exists, is a point that is common to both functions / graphs.

OpenStudy (ikram002p):

check the last one its \(y=x^2\) \(y=2x +9\) so \(x^2=2x+9\) is it true ?

OpenStudy (anonymous):

no

OpenStudy (ikram002p):

so what do u think nw ? what the answer might be ?

OpenStudy (anonymous):

i dont know, im a little confused now

OpenStudy (mathmale):

Please hold on for a moment. I don't think it appropriate to ask, "is it true," in regard to the equation\[x^2=2x+9.\]This equation is not (and was not supposed to be) an identity, but rather a conditional equality. "conditional" in this case means taht the equation is true only for certain x values. Your job is to determine what those values are.

OpenStudy (ikram002p):

^_^

OpenStudy (mathmale):

Here's a quick summary: x2 = x2 + 2x + 9 is given. For now, forget about "setting this up as a system of equations." Subtracting x^2 from both sides, 0=2x+9, or 2x=-9, or x= ???

OpenStudy (mathmale):

In other words, solve 2x=-9 for x.

OpenStudy (mathmale):

Sorry, I need to go elsewhere. Glad to be of help, but I do hope for reasonably quick responses to my posts. Perhaps later?